OpenStudy (anonymous):

If x = 25 units, y = 30 units, z = 6 units, and h = 20 units, then what is the surface area of the triangular prism shown above?

OpenStudy (ltrout):

Do you know the formula for the surface area?

OpenStudy (ltrout):

The formula is SA = wh + lw + lh + ls. Using the information you provided: w = 25 l = 30 s = 6 h = 20 Back to the equation: Sa = (25)(20) + (30)(25) + (30)(20) + (30)(6) = ? Can you solve from here?

OpenStudy (welshfella):

where is the diagram?

OpenStudy (ltrout):

@welshfella I realized that too. But we shouldn't need it since he gave us the sizes?

OpenStudy (welshfella):

its a triangular prism so what exactly does x , y an z mean?

OpenStudy (ltrout):

x, y, and z would be the length, width, height I'm assuming.

OpenStudy (welshfella):

what is h then?

OpenStudy (ltrout):

height

OpenStudy (welshfella):

height of what?

OpenStudy (welshfella):

you said z = height also

OpenStudy (welshfella):

anyway - i guess we'll never know as he's disappeared...

OpenStudy (ltrout):

Sorry about that. I'm assuming that x, y, z = length, width, and a side? And then h = height. Could be wrong, but you're right, he's gone.

OpenStudy (welshfella):

z could be the height of the triangle and h = height of the prism but without a diagram we can't be sure